help please!!!!
A banana peel weighs 1/8 the total weight of a banana. If an unpeeled banana balances a peeled banana of the same weight plus 7/8 of an ounce. How much does the banana weigh with peel?

Answer by josmiceli(19441)

I'll call the peeled banana fruit
fruit + peel = banana
The 2nd part of the problem says
banana = fruit + 7/8 (in ounces)
The 1st part says
peel = (1/8)*banana
You know that
banana = fruit + peel
so, peel = 7/8 ounce
7/8 = (1/8)*banana
multiply both sides by 8
banana = 7 ounces answer
also note that
peel / banana = 1/8
(7/8) / 7 = 1/8
multiply both sides by 7
7/8 = 7/8